OpenStudy (anonymous):

Find the perpendicular slope of the line 3X-9Y=15

OpenStudy (anonymous):

Solve for y y = (3x -15)/9 = (1/3)x - 5/3 The slope is 1/3. The line perpendicular to this has slope -1/(1/3) = -3